3rd RESCEU Symposium ‘Particle Cosmology’ Program(Poster)

1) Asada, H.  Gravitational Lensing in the Clumpy Universe and Cosmological Parameters
2) Fujisaki, H.  Thermal Cosmological Constant in the Heterotic String Thermofield Dynamics
3) Fukui, T.  The Relation of the Generalized Scalar-Tensor Theory

with the 5D Space-Time-Mass Theory

4) Gen, U.  Timelike infinity and Asymptotic Symmetry
5) Hara, T.  On the Formation of Astronomical Objects in the Cosmic String Scheme
6) Ishibashi, A.  Anisotropy in Light Propagation over Cosmological Distance
7) Kanazawa, T.  Hybrid Inflation and Axionic Isocurvature Density Perturbation in Supergravity
8) Mukohyama, S.  Thermodynamics of Entanglement
9) Nakamura, T. T.  The spectrum of cosmological perturbations produced by multi-component

inflaton to second order in the slow-roll approximation

10) Nagasawa, M.  Baryon Number Violation through Interaction between Sphaleron

and Cosmological Defect

11) Nagataki, S.  Explosive Nucleosynthesis and Matter Mixing

in Axisymmetric Type II Supernova Explosion

12) Nakazawa, N.  Proper Time Interpretation of Stochastic Time and 4D Lattice Quantum Gravity
13) Ohkuwa, Y.  On Third Quantization of Kaluza-Klein Cosmology
14) Orito, M.  Inhomogeneous Big Bang Nucleosynthesis and Lepton Asymmetry
15) Sakai, N.  Inflating Monopoles in Brans-Dicke Theory
16) Soda, J.  Quantum Back Reaction and the Non-Singular Cosmology in 2-Dimensions
17) Taruya, A.  Cosmological Perturbation with Two Scalar Fields in Reheating after Inflation
18) Tsubaki, S.  Cosmic Rays and Gamma Ray in Clusters of Galaxies
19) Woon, S. C.  Analytic Continuation of Quantum Operators and Imaginary Space

vs. Imaginary Time Formulation

20) Yamaguchi, M.  Probability Distribution Function for the Coarse-grained Field